# Direct Bias-Correction Term Estimation for Propensity Scores and Average Treatment Effect Estimation

### v1

@@ 7 Simulation studies
We assess the empirical performance of our method through simulation studies, evaluating ATE estimation error across a range of scenarios. We compare our approach with CBPS (Imai & Ratkovic, 2013) and RieszNet (Chernozhukov et al., 2022a). Because our least squares is equivalent to Resz regression, we include RieszNet primarily as a numerical check of equivalence, noting architectural differences.
@@
For propensity score estimation, we employ a three-layer neural network with an Exponential Linear Unit (ELU) activation function for each hidden layer ($100$ nodes per layer). The final output layer applies a sigmoid function to ensure that the estimated propensity scores remain in $(0,1)$. We use this model for our method, logistic regression, and CBPS. For RieszNet, we adopt the DragonNet architecture proposed in Shi et al. (2019), following the original implementation by Chernozhukov et al. (2022a). For each propensity score estimation method, including ours, we compute both the IPW and AIPW estimators using the estimated scores. Additionally, we include the direct method (DM) estimator with neural networks for comparison. In each case, the expected conditional outcomes are estimated using a three-layer neural network (100 nodes per hidden layer, with ELU activation). As a baseline, we also consider the DM estimator with linear models.
@@ Appendix A Density-Ratio Estimation (DRE)
Appendix B Bregman divergence
Here, we show the equivalence between
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big),$
where $\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)\coloneqq\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$, and
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}_{g}\big(r\big),$
where $\mathrm{BR}_{g}\big(r\big)=\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\Big(-g(r(d,X))+\partial g(r(d,X))r(d,X)\Big)-\partial g(r(d,X))\Big]$.
This can be shown as follows:
$r^{*}=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\mathrm{BR}^{\dagger}_{g}\big(r_{0}\mid r\big)$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(g(r_{0}(d,X))-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))\Big(r_{0}(d,X)-r(d,X)\Big)\right)\Big]$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))r(d,X)\right)\Big]-{\mathbb{E}}\Big[\mathbbm{1}[D=d]\partial g(r(d,x))r_{0}(d,X)\Big]$
$=\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\sum_{d\in\{1,0\}}{\mathbb{E}}\Big[\mathbbm{1}[D=d]\left(-g(r(d,X))-\partial g(r(d,x))r(d,X)\right)\Big]-{\mathbb{E}}\Big[\partial g(r(d,x))\Big].$
Here, we used
${\mathbb{E}}[\mathbbm{1}[D=1]r_{0}(1,X)\mid X]={\mathbb{E}}[e_{0}(X)r_{0}(1,X)\mid X]=1.$
We can show the same result for the case with $D=0$.
@@ Appendix C Estimation of the average treatment effect for the treated (ATT)
$\widetilde{h}_{0}(D,X,e_{0},\pi_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{e_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}(1-e_{0}(X))},$
@@
Let $w_{0}(x)\coloneqq\frac{e_{0}(X)\mathbbm{1}[D=0]}{(1-e_{0}(X))}$. Then, we denote the bias-correction term as
$\widetilde{h}_{0}(D,X,w_{0},\pi_{0})\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{w_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}}.$
@@
$w^{*}=\operatorname*{arg\,min}_{r\in\mathcal{R}}\left\{-2{\mathbb{E}}_{1}\left[w(X)\right]+{\mathbb{E}}\left[w(X)^{2}\mathbbm{1}[D=0]\right]\right\},$
@@ D.1 Rademacher complexity
$\mathfrak{R}_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}\sigma_{i}f(X_{i}).$
@@ D.2 Local Rademacher complexity bound
$Pf\coloneqq{\mathbb{E}}[f(X)],$
$P_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(X_{i}).$
@@
Let $\mathcal{F}$} be a class of functions that map $\mathcal{X}$} into $[a,b]$. Assume that there is some $r>0$ such that for every $f\in{\mathcal{F}}$, $\mathrm{Var}(f(X))\leq r$. Then, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that
@@ Appendix E Basic inequalities
$\widehat{r}\coloneqq\operatorname*{arg\,min}_{r\in{\mathcal{R}}}\widehat{{\mathcal{L}}}_{n}(r)+\lambda J(r),$
where $J(r)$ is some regularization term.
Throughout the proof, we use the following basic inequalities that hold for $\widehat{r}$.
@@
[…]
$\leq\frac{1}{n}\sum^{n}_{i=1}\Big(-2r_{0}(1,X_{i})-2r_{0}(0,X_{i})+\mathbbm{1}[D_{i}=1]r_{0}(1,X_{i})^{2}+\mathbbm{1}[D_{i}=0]r_{0}(0,X_{i})^{2}\Big)+\lambda J(r_{0}).$
For a function $f\colon{\mathcal{X}}\to{\mathbb{R}}$ and $X$ following $P$, let us denote the sample mean and expectation as
$\int f{\mathrm{d}}P_{n}$ $\coloneqq\int f(\cdot){\mathrm{d}}P_{n}\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(X_{i}),$
$\int f{\mathrm{d}}P$ $\coloneqq\int f(\cdot){\mathrm{d}}P\coloneqq{\mathbb{E}}[f(X)].$
Following this notation, we also express this inequality as
$\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{n}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ +\int\mathbbm{1}[D_{i}=1]\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}+\int\mathbbm{1}[D_{i}=0]\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,X)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ +\lambda J(\widehat{r})-\lambda J(r_{0})\leq 0.$
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
$\leq\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
$\ \ \ +\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid X)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$
@@
Let $e_{0}(1\mid x)=e_{0}(x)$ and $e_{0}(0\mid x)=1-e_{0}(x)$. Then, the following holds for the $L_{2}$ estimation error:
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
$=\left\|\mathbbm{1}[D=1]\big(\widehat{r}(1,X)-r_{0}(1,X)\big)-\mathbbm{1}[D=0]\big(\widehat{r}(0,X)-r_{0}(0,X)\big)\right\|^{2}_{P_{0}}$
$=\left\|e_{0}(1\mid\cdot)\big(\widehat{r}(1,X)-r_{0}(1,X)\big)\right\|^{2}_{P_{0}}+\left\|e_{0}(1\mid 0)\big(\widehat{r}(0,X)-r_{0}(0,X)\big)\right\|^{2}_{P_{0}}$
$=\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}.$
@@
$\left\|h(D,X;\widehat{r})-h(D,X;r_{0})\right\|^{2}_{P_{0}}$
$\leq\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
$\ \ \ -\int-2\big(\widehat{r}_{n}(1,\cdot)-r^{*}(1,\cdot)\big){\mathrm{d}}P_{n}-\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ -\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r^{*}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}-\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r^{*}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ -\lambda J(\widehat{r})+\lambda J(r_{0})$
$=\int-2\big(\widehat{r}_{n}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{0}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r_{0}(0,\cdot)\big){\mathrm{d}}P_{0}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{0}+\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{0}$
$\ \ \ -\int-2\big(\widehat{r}_{n}(1,\cdot)-r^{*}(1,\cdot)\big){\mathrm{d}}P_{n}-\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ -\int e_{0}(1\mid\cdot)\Big(\widehat{r}_{n}(1,\cdot)^{2}-r^{*}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}-\int e_{0}(0,\cdot)\Big(\widehat{r}_{n}(0,\cdot)^{2}-r^{*}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ +\int-2\big(r_{0}(1,\cdot)-r_{0}(1,\cdot)\big){\mathrm{d}}P_{n}+\int-2\big(\widehat{r}_{n}(0,\cdot)-r^{*}(0,\cdot)\big){\mathrm{d}}P_{n}$
$\ \ \ +\int e_{0}(1\mid\cdot)\Big(r_{0}(1,\cdot)^{2}-r_{0}(1,\cdot)^{2}\Big){\mathrm{d}}P_{n}+\int e_{0}(0,\cdot)\Big(r_{0}(0,\cdot)^{2}-r_{0}(0,\cdot)^{2}\Big){\mathrm{d}}P_{n}$
$\ \ \ -\lambda J(\widehat{r})+\lambda J(r_{0}).$
Here, we used
$\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}$ $=2\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}-\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}$
$=2\int r_{0}(1,\cdot){\mathrm{d}}P_{0}-\int\mathbbm{1}[d=1]r_{0}(1,\cdot)^{2}{\mathrm{d}}P_{0}.$
@@
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})$ (2)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg)$
$\ \ \ +\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)$ (3)
in Proposition E.2.
Here, since $r_{0}\in{\mathcal{H}}$, it holds that $r^{*}=r_{0}$, which implies that
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(r^{*}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}P_{n}+\int e_{0}(d\mid\cdot)\Big(r^{*}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}P_{n}\Bigg)=0.$
Therefore, we consider bounding the first sum
$\sum_{d\in\{1,0\}}\Bigg(\int-2\big(\widehat{r}_{n}(d,\cdot)-r_{0}(d,\cdot)\big){\mathrm{d}}(P_{0}-P_{n})+\int e_{0}(d,\cdot)\Big(\widehat{r}_{n}(d,\cdot)^{2}-r_{0}(d,\cdot)^{2}\Big){\mathrm{d}}(P_{0}-P_{n})\Bigg).$
We can bound this term by using the empirical-process arguments.

### v2

@@ 6 Simulation studies
We assess the performance of our method through simulation studies, evaluating ATE estimation error. We denote our direct bias-correction term estimation methods as DBC (LS) when using the squared loss, and DBC (TL) when using the tailored loss. We compare our approach with ATE estimators using propensity score estimated by maximum likelihood estimation (MLE), CBPS (Imai & Ratkovic, 2013a), and RieszNet (Chernozhukov et al., 2022a). Because our DBC (LS) is equivalent to Resz regression, we include RieszNet primarily as a numerical check of equivalence, noting architectural differences. In this section, for simplicity, we do not apply cross-fitting. We also conduct experiments in Appendices I and J using synthetic and semi-synthetic data, respectively, in which we apply cross-fitting.
@@
We model $h_{0}$ by modeling $e_{0}$. To model $e_{0}$, we use a three-layer neural network with an Exponential Linear Unit (ELU) activation function for each hidden layer ($100$ nodes per layer). The final output layer applies a sigmoid function to ensure that the estimated propensity scores remain in $(0,1)$. We use this model for our method, logistic regression, and CBPS. For RieszNet, we adopt the DragonNet architecture proposed in Shi et al. (2019), following Chernozhukov et al. (2022b). For each method, including ours, we compute both the IPW and AIPW estimators using the estimated scores. Additionally, we include the direct method (DM) estimator with neural networks for comparison. In each case, the expected conditional outcomes are estimated using a three-layer neural network (100 nodes per hidden layer, with ELU activation). As a baseline, we also consider the DM estimator with linear models.
@@ Appendix A Density-Ratio Estimation (DRE)
Appendix B Silverman’s trick
Note that minimization of the Bregman divergence with the KL divergence loss is equal to
@@ Appendix C Estimation of the average treatment effect for the treated (ATT)
$\widetilde{h}_{0}(D,X)\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{e_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}(1-e_{0}(X))},$
@@
Let $w_{0}(x)\coloneqq\frac{e_{0}(X)}{(1-e_{0}(X))}$. Then, we denote the bias-correction term as
$\widetilde{h}_{0}(D,X)\coloneqq\frac{\mathbbm{1}[D=1]}{\pi_{0}}-\frac{w_{0}(X)\mathbbm{1}[D=0]}{\pi_{0}}.$
@@
$w^{*}=\operatorname*{arg\,min}_{r\in\mathcal{R}}\left\{-2{\mathbb{E}}\left[w(X)\right]+{\mathbb{E}}\left[w(X)^{2}\mathbbm{1}[D=0]\right]\right\},$
@@ D.1 Rademacher complexity
$\mathfrak{R}_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}\sigma_{i}f(W_{i}).$
@@ D.2 Local Rademacher complexity bound
$Pf\coloneqq{\mathbb{E}}[f(W)],$
$P_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(W_{i}).$
@@
Let $\mathcal{F}$} be a class of functions that map $\mathcal{X}$} into $[a,b]$. Assume that there is some $r>0$ such that for every $f\in{\mathcal{F}}$, $\mathrm{Var}(f(W))\leq r$. Then, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that
@@ Appendix E Basic inequalities
E.1 Strong convexity
Lemma E.1 ($L_{2}$ distance bound from Lemma 4 in Kato & Teshima (2021)).
If $\inf_{h\in(-\infty),\infty}g^{\prime\prime}(h)>0$, then there exists $\mu>0$ such that for all $h\in{\mathcal{H}}$,
$\|h-h_{0}\|^{2}_{2}\leq\frac{2}{\mu}\Big(\mathrm{BR}_{g}(h)-\mathrm{BR}_{g}(h_{0})\Big)$
holds.
From the strong convexity and Lemma E.1, we have
$\frac{\mu}{2}\|\widehat{h}_{n}-h_{0}\|^{2}_{2}\leq\mathrm{BR}_{g}(\widehat{h}_{n})-\mathrm{BR}_{g}\big(h_{0}\big).$
@@ E.1 Strong convexity
$\widehat{h}\coloneqq\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\widehat{{\mathcal{L}}}_{n}(h)+\lambda J(h),$
where $J\left(h\right)$ is some regularization term.
E.2 Preliminary
@@
$\widehat{\mathrm{BR}}_{g}(\widehat{h})+\lambda J(\widehat{h})\leq\widehat{\mathrm{BR}}_{g}(h^{*})+\lambda J(h^{*}),$
where recall that
$\widehat{\mathrm{BR}}_{g}(h)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(-g(h(D_{i},X_{i}))+\partial g(h(D_{i},X_{i}))h(D_{i},X_{i})-\partial g(h(1,X_{i}))-\partial g(h(0,X_{i}))\Big).$
Let $Z\in{\mathcal{Z}}$ be a random variable with a space $\mathcal{Z}$}, and $\{Z_{i}\}^{n}_{i=1}$ be its realizations. For a function $f\colon{\mathcal{Z}}\to{\mathbb{R}}$ and $X$ following $P$, let us denote the sample mean as
$\widehat{{\mathbb{E}}}[f(Z)]\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(Z_{i}).$
We also denote $\widehat{{\mathbb{E}}}[f(Z)]-{\mathbb{E}}[f(Z)]=(\widehat{{\mathbb{E}}}-{\mathbb{E}})f(Z)$
E.3 Risk bound
Recall that
$\widehat{\mathrm{BR}}_{g}(h)=\frac{1}{n}\sum^{n}_{i=1}\Big(-g(h(D_{i},X_{i}))+\partial g(h(D_{i},X_{i}))h(D_{i},X_{i})-\partial g(h(1,X_{i}))-\partial g(h(0,X_{i}))\Big).$
Let us define
$L(h,D,X)\coloneqq-g(h(D,X))+\partial g(h(D,X))h(D,X)-\partial g(h(1,X))-\partial g(h(0,X)),$
and we can write
$\widehat{\mathrm{BR}}_{g}(h)=\widehat{{\mathbb{E}}}\big[L(h,D,X)\big]$
Then, from Proposition E.2, we have
$\widehat{{\mathbb{E}}}\big[L(h^{*},D,X)\big]-\widehat{{\mathbb{E}}}\big[L(\widehat{h}_{n},D,X)\big]+\lambda J(\widehat{h})-\lambda J(h^{*})\geq 0.$
Throughout the proof, we use the following basic inequalities that hold for $\widehat{h}$.
@@ E.3 Risk bound
$\frac{\mu}{2}\left\|\widehat{h}_{n}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}$
@@
From the strong convexity and Lemma E.1, we have
$\frac{\mu}{2}\|\widehat{h}_{n}-h_{0}\|^{2}_{2}\leq\mathrm{BR}_{g}(\widehat{h}_{n})-\mathrm{BR}_{g}\big(h_{0}\big)={\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right].$
@@
$\frac{\mu}{2}\left\|\widehat{h}(D,X)-h_{0}(D,X)\right\|^{2}_{L_{2}(P_{0})}$
$\leq{\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$={\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$\ \ \ \ \ \ -\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$\ \ \ \ \ \ +\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$\leq{\mathbb{E}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$\ \ \ \ \ \ -\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$\ \ \ \ \ \ +\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right]$
$\ \ \ \ \ \ -\widehat{{\mathbb{E}}}\left[L(\widehat{h}_{n},D,X)-L(h^{*},D,X)\right]+\lambda J(\widehat{h})-\lambda J(h_{0}).$
@@
$\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{h}_{n},D,X)-L(h_{0},D,X)\right],$ (1)
in Proposition E.3. We can bound this term by using the empirical-process arguments.
Note that since $h_{0}\in{\mathcal{H}}$, it holds that $h^{*}=h_{0}$, which implies that